1. The probability that Tim will sink a foul shot is 70%. If Tim attempts 10


   foul shots, what is the probability that

a)    he sinks exactly 7 shots

b)   he sinks at least 7 shots

c)    he sinks at most 7 shots

d)   he sinks between 4 and 6 shots, inclusive.





1
Expert's answer
2022-06-27T09:44:14-0400

Let "X=" the number of shots sinked by Tim: "X\sim Bin (n, p)."

Given "n=10, p=0.7, q=1-p=0.3."

a)


"P(X=7)=\dbinom{10}{7}(0.7)^{7}(0.3)^{10-7}=0.266827932"

b)


"P(X\ge 7)=P(X=7)+P(X=8)+P(X=9)"

"+P(X=10)=\dbinom{10}{7}(0.7)^{7}(0.3)^{10-7}"

"+\dbinom{10}{8}(0.7)^{8}(0.3)^{10-8}+\dbinom{10}{9}(0.7)^{9}(0.3)^{10-9}"

"+\dbinom{10}{10}(0.7)^{10}(0.3)^{10-10}=0.6496107184"

c)


"P(X\le 7)=1-P(X=8)-P(X=9)"

"-P(X=10)=1-\dbinom{10}{8}(0.7)^{8}(0.3)^{10-8}"

"-\dbinom{10}{9}(0.7)^{9}(0.3)^{10-9}-\dbinom{10}{10}(0.7)^{10}(0.3)^{10-10}"

"=0.6172172136"

d)


"P(4\le X\le 6)=P(X=4)+P(X=5)"

"+P(X=6)=\dbinom{10}{4}(0.7)^{4}(0.3)^{10-4}"

"+\dbinom{10}{5}(0.7)^{5}(0.3)^{10-5}+\dbinom{10}{6}(0.7)^{6}(0.3)^{10-6}"

"=0.3397972032"